# From Path Coefficients to Targeted Estimands: A Comparison of Structural Equation Models (SEM) and Targeted Maximum Likelihood Estimation (TMLE)

### v1

@@ 3.2.1 Estimating NDE and NIE via SEM
A commonly used approach for simply the conditional expectation in identifying NIE and NDE is encoding it as a parametric SEM[40]. Suppose our continuous mediator and continuous outcomes are specified under the following structural equations
@@
where $U$ is the unobserved and uncorrelated exogenous variables, and $g_{A}$ is some known function. We further assume that $M$ has a probability density function $f_{M}$ with respect to some dominating measure and discrete $\boldsymbol{W}$ takes values in some state space $\mathcal{W}\subset\mathbb{R}^{d}$. Since the parameteric assumptions are specified here, the direct and indirect effects can be derived in a closed form [46, 41].
@@
This shows that under correct model specification, estimating the NIE and NDE is equivalent to estimating the path coefficients in SEM. We may also identify it using the path analysis and coefficients; there are two paths from $A$ to $Y$: $A\rightarrow M\rightarrow Y$ and $A\rightarrow Y$, where the former passes $M$, which would be the indirect effect and the latter would be the direct effect. Given the independence of $U_{M}$ and $U_{Y}$, $Y$ would follow a normal distribution with the model-induced distribution parameters. Hence, the parameter estimation could be done by maximizing the model-induced likelihood with the following Wald test for inference. For NIE, the inference is commonly done via the Delta method or via Bootstrap [59].
@@ 3.2.2 Estimating NDE and NIE via TMLE
The parametric structural equations we specified simplify the estimation of the NIE and NDE, but they may be restrictive and might not accurately reflect the true data-generating mechanism. TMLE relaxes these linear assumptions and instead identifies the causal effects based on the NPSEM. We give the steps for estimating NDE via TMLE introduced by [68], while the implementation for NIE can be derived as an analog. Without assuming the parametric assumptions, we instead arrive at the following NPSEM
@@
which solves the EIF and attains asymptotic efficiency under regularity conditions, including consistency of the nuisance estimators for $\bar{Q}_{Y}$, $g$, and $Q_{M}$, achieving the semiparametric efficiency bound $\mathrm{Var}(D^{*}(P_{0}))$. For software availability, [47] provides a R package called medoutcon for estimating NIE and NDE using TMLE with the super learners implemented.
@@ 4.0.1 Methods for Average Treatment Effects
We consider a similar data-generating process in [53], which is equivalent to the following DAG shown in Figure 4.
@@
Figure 5: A structural causal graph adapted from [53] illustrating the relationships between exogenous variables $W_{1}-W_{4}$, binary treatment $A$ and the outcome variable $Y$.
@@
$Y|A,\boldsymbol{W}\sim\mathcal{N}\left(-1+\psi\times A+0.1\times W_{1}+0.35\times W_{2}+0.25\times W_{3}+0.2\times W_{4}+3.0\times W_{2}\times W_{4},1\right),$
@@
$Y|A,\boldsymbol{W}\sim\mathcal{N}\left(-1+\psi A+0.1\cdot W_{1}+0.35\cdot W_{2}+0.25\cdot W_{3}^{4}+0.2\cdot W_{4}^{4}+3.0\cdot W_{2}\cdot W_{4},1\right),$
@@
We conducted a Monte Carlo simulation with $n_{\text{sim}}=1000$ with various sample size $n\in\{200,500,750,1000,1500,2000,3000\}$ using both lm and tmle[43] in R and assess the consistency and efficiency using (i) Relative Bias: $n_{\text{sim}}^{-1}\sum_{i=1}^{n_{\text{sim}}}\frac{\hat{\psi}-\psi}{\psi}$; (ii) $95\%$-CI coverage: $n_{\text{sim}}^{-1}\sum_{i=1}^{n}\mathbf{1}_{\psi\in(\ell,u)}$, where $\ell,u$ denote the lower and upper bounds of $95\%$-CI; (iii) Statistical Power against $\psi=0$: $n_{\text{sim}}^{-1}\sum_{i=1}^{n_{\text{sim}}}\mathbf{1}_{0\not\in(\ell,u)}$; (iv) Standardized RMSE: $\sqrt{n_{\text{sim}}^{-1}\sum_{i=1}^{n_{\text{sim}}}\left(\frac{\hat{\psi}-\psi}{\psi}\right)^{2}}$. Instead of linear regression, one could also use the SEM with a logit link to estimate the path coefficient. The superlearner libraries used for tmle are Generalized Linear Model(glm, glm.interaction); generalized additive model(gam); random forest(ranger) and eXtreme Gradient Boosting(xgboost). We then repeat the experiment for CATE, by adding an interaction term between $A$ and $W_{1}$
$Y|A,\boldsymbol{W}\sim\mathcal{N}\left(-1+\psi A+0.5A\times W_{1}+0.1\times W_{1}+0.35\times W_{2}+0.25\times W_{3}+0.2\times W_{4}+3.0\cdot W_{2}\times W_{4},1\right),$
@@ 4.0.2 Methods for Mediation Analysis
The mediation effects under this data-generating process could be identified to be $\psi_{\text{NIE}}=1$ and $\psi_{\text{NDE}}=2$. To conduct the mediation analysis, both SEM and TMLE are fitted and assessed. SEM was implemented using lavaan in R, and the confidence interval with the corresponding inference were based on bootstrap with a number of samples equal to 1000. The TMLE for the mediation analysis was implemented using medoutcon [47] with Generalized Linear Models and Random Forest as the superlearner algorithms being used to obtain the estimates in each of the steps. Due to the computational costs, we conduct a Monte Carlo simulation with $n_{\text{sim}}=200$ with sample sizes $n\in\{500,800,1000,1500,2000,2500,3000,5000\}$ and assess both of the estimators using the same metrics that were computed for ATE and CATE.
@@ 5.1 Data and Study Design
The analysis draws from six waves between 2012 and 2022 of the China Family Panel Study (CFPS), a nationally representative longitudinal survey. The sample consists of $2881$ rural children aged from six to fifteen in the 2012 wave and newly sampled children aged from eight to fifteen in the 2014 wave, tracked for high school enrollment at ages 15 or 16 in subsequent waves. The binary outcome is high school enrollment[50]. Following [62], multidimensional poverty was measured using nine indicators across three dimensions: health (nutrition, child mortality), education (parental years of schooling, children’s elementary school attendance), and living standards (cooking fuel, sanitation, drinking water, electricity, housing assets). Each dimension and its indicators within each dimension were equally weighted. A child was considered multidimensionally poor if their weighted deprivation score was 0.33 or higher. To satisfy the unconfoundedness assumption, covariates include individual factors (age, gender, siblings), family characteristics (household registration, living arrangements, parental education and occupation, political identity, $\log$-per capita income, Dibao receipt), school quality, and province fixed effects. All the above covariates were drawn from the first wave for each sample. Additionally, we selected children’s educational expenditure as a mediator, measured in two ways: first, as the logarithm of the absolute expenditure amount; second, as the proportion of total household income spent on education.
@@ 5.2 Estimation of (Conditional)Average Treatment Effects
To assess the average impact of multidimensional poverty on children’s access to high school, parametric logistic regression using glm in R, and TMLE with default superlearners in tmle, including Regularized GLM and Bayesian Additive Regression Trees [43] were implemented.
@@
This aligns with the previous research, which showed in the education stage, boys are often more sensitive to familial disadvantages [35, 54, 51]. Families often impose higher educational expectations on boys, reflecting broader societal norms that link male identity to achievement. In the context of poverty, however, limited economic resources and constrained parental support can impede boys’ capacity to meet these expectations, exacerbating vulnerabilities and leading to lower academic attainment, diminished motivation, and increased risk of adverse educational outcomes relative to girls.
@@ 5.3 Mediation Analysis
Beyond estimating (conditional) average treatment effects, sociologists are often equally concerned with the mechanisms through which causal relationships unfold. To assess the direct and indirect effects, SEM using lavaan [60] and TMLE using medoutcon [47] in R were implemented. To account for the potential non-linearity and sparsity, algorithms including Random Forest, LASSO, Gradient Boosting, and Generalized Additive Model are specified in the superlearner when implementing the TMLE.

### v2

@@ 3.2.1 Estimating NDE and NIE via SEM
A commonly used approach for simplifying the conditional expectation in identifying NIE and NDE is encoding it as a parametric SEM[de-stavola-2014]. Suppose our continuous mediator and continuous outcomes are specified under the following structural equations
@@
where $U$ is the unobserved and uncorrelated exogenous variables, and $g_{A}$ is some known function. We further assume that $M$ has a probability density function $f_{M}$ with respect to some dominating measure and discrete $\boldsymbol{W}$ takes values in some state space $\mathcal{W}\subset\mathbb{R}^{d}$. Since the parameteric assumptions are specified here, the direct and indirect effects can be derived in a closed form [gunzler-2013, ding-2024].
@@
This shows that under correct model specification, estimating the NIE and NDE is equivalent to estimating the path coefficients in SEM. We may also identify it using the path analysis and coefficients; there are two paths from $A$ to $Y$: $A\rightarrow M\rightarrow Y$ and $A\rightarrow Y$, where the former passes $M$, which would be the indirect effect and the latter would be the direct effect. Given the independence of $U_{M}$ and $U_{Y}$, $Y$ would follow a normal distribution with the model-induced distribution parameters. Hence, the parameter estimation could be done by maximizing the model-induced likelihood with the following Wald test for inference. For NIE, the inference is commonly done via the Delta method or via Bootstrap [rosseel-2012].
@@ 3.2.2 Estimating NDE and NIE via TMLE
The parametric structural equations we specified simplify the estimation of the NIE and NDE, but they may be restrictive and might not accurately reflect the true data-generating mechanism. TMLE relaxes these linear assumptions and instead identifies the causal effects based on the NPSEM. We give the steps for estimating NDE via TMLE introduced by [zheng-2012], while the implementation for NIE can be derived as an analog. Without assuming the parametric assumptions, we instead arrive at the following NPSEM
@@
which solves the EIF and attains asymptotic efficiency under regularity conditions, including consistency of the nuisance estimators for $\bar{Q}_{Y}$, $g$, and $Q_{M}$, achieving the semiparametric efficiency bound $\mathrm{Var}(D^{*}(P_{0}))$. For software availability, [hejazi-2022] provides a R package called medoutcon for estimating NIE and NDE using TMLE with the super learners implemented.
@@ 4.0.1 Methods for Average Treatment Effects
We consider a similar data-generating process in [luquefernandez-2018], which is equivalent to the following DAG shown in Figure 5.
@@
Figure 5: A structural causal graph adapted from [luquefernandez-2018] illustrating the relationships between exogenous variables $W_{1}-W_{4}$, binary treatment $A$ and the outcome variable $Y$.
@@
$Y|A,\boldsymbol{W}\sim\mathcal{N}\left(-1+\psi\times A+0.1\times W_{1}+0.35\times W_{2}+0.25\times W_{3}+0.2\times W_{4}+3.0\times W_{2}W_{4},1\right),$
@@
$Y|A,\boldsymbol{W}\sim\mathcal{N}\left(-1+\psi\times A+0.1\times W_{1}+0.35\times W_{2}+0.25\times W_{3}^{4}+0.2\times W_{4}^{4}+3.0\times W_{2}W_{4},1\right),$
@@
We conducted a Monte Carlo simulation with $n_{\text{sim}}=1000$ with various sample size $n\in\{200,500,750,1000,1500,2000,3000\}$ using both lm and tmle[gruber-2012] in R and assess the consistency and efficiency using (i) Relative Bias: $n_{\text{sim}}^{-1}\sum_{i=1}^{n_{\text{sim}}}\frac{\hat{\psi}-\psi}{\psi}$; (ii) $95\%$-CI coverage: $n_{\text{sim}}^{-1}\sum_{i=1}^{n}\mathbf{1}_{\psi\in(\ell,u)}$, where $\ell,u$ denote the lower and upper bounds of $95\%$-CI; (iii) Statistical Power against $\psi=0$: $n_{\text{sim}}^{-1}\sum_{i=1}^{n_{\text{sim}}}\mathbf{1}_{0\not\in(\ell,u)}$; (iv) Standardized RMSE: $\sqrt{n_{\text{sim}}^{-1}\sum_{i=1}^{n_{\text{sim}}}\left(\frac{\hat{\psi}-\psi}{\psi}\right)^{2}}$. Instead of linear regression, one could also use the SEM with a logit link to estimate the path coefficient. The superlearner libraries used for tmle are generalized linear model(glm, glm.interaction); generalized additive model(gam); random forest(ranger) and eXtreme Gradient Boosting(xgboost). We then repeat the experiment for CATE, by adding an interaction term between $A$ and $W_{1}$
$Y|A,\boldsymbol{W}\sim\mathcal{N}\left(-1+\psi A+0.5\times AW_{1}+0.1\times W_{1}+0.35\times W_{2}+0.25\times W_{3}+0.2\times W_{4}+3.0\times W_{2}W_{4},1\right),$
@@ 4.0.2 Methods for Mediation Analysis
The mediation effects under this data-generating process could be identified to be $\psi_{\text{NIE}}=1$ and $\psi_{\text{NDE}}=2$. To conduct the mediation analysis, both SEM and TMLE are fitted and assessed. SEM was implemented using lavaan in R, and the confidence interval with the corresponding inference were based on bootstrap with a number of samples equal to 1000. The TMLE for the mediation analysis was implemented using medoutcon [hejazi-2022] with Generalized Linear Models and Random Forest as the superlearner algorithms being used to obtain the estimates in each of the steps. Due to the computational costs, we conduct a Monte Carlo simulation with $n_{\text{sim}}=200$ with sample sizes $n\in\{500,800,1000,1500,2000,2500,3000,5000\}$ and assess both of the estimators using the same metrics that were computed for ATE and CATE.
@@ 5.1 Data and Study Design
The analysis draws from six waves between 2012 and 2022 of the China Family Panel Study (CFPS), a nationally representative longitudinal survey. The sample consists of $2881$ rural children aged from six to fifteen in the 2012 wave and newly sampled children aged from eight to fifteen in the 2014 wave, tracked for high school enrollment at ages 15 or 16 in subsequent waves. The binary outcome is high school enrollment [lei-2021]. Following [szaflarski-2019], multidimensional poverty was measured using nine indicators across three dimensions: health (nutrition, child mortality), education (parental years of schooling, children’s elementary school attendance), and living standards (cooking fuel, sanitation, drinking water, electricity, housing assets). Each dimension and its indicators within each dimension were equally weighted. A child was considered multidimensionally poor if their weighted deprivation score was 0.33 or higher. To satisfy the unconfoundedness assumption, covariates include individual factors (age, gender, siblings), family characteristics (household registration, living arrangements, parental education and occupation, political identity, $\log$-per capita income, Dibao receipt), school quality, and province fixed effects. All the above covariates were drawn from the first wave for each sample. Additionally, we selected children’s educational expenditure as a mediator, measured in two ways: first, as the logarithm of the absolute expenditure amount; second, as the proportion of total household income spent on education.
@@ 5.2 Estimation of (Conditional)Average Treatment Effects
To assess the average impact of multidimensional poverty on children’s access to high school, parametric logistic regression using glm in R, and TMLE with default superlearners in tmle, including Regularized GLM and Bayesian Additive Regression Trees [gruber-2012] were implemented.
@@
This aligns with the previous research, which showed in the education stage, boys are often more sensitive to familial disadvantages [autor-2019, owens-2016, lei-2020]. Families often impose higher educational expectations on boys, reflecting broader societal norms that link male identity to achievement. In the context of poverty, however, limited economic resources and constrained parental support can impede boys’ capacity to meet these expectations, exacerbating vulnerabilities and leading to lower academic attainment, diminished motivation, and increased risk of adverse educational outcomes relative to girls.
@@ 5.3 Mediation Analysis
Beyond estimating (conditional) average treatment effects, sociologists are often equally concerned with the mechanisms through which causal relationships unfold. To assess the direct and indirect effects, SEM using lavaan [schumacker-2015] and TMLE using medoutcon [hejazi-2022] in R were implemented. To account for the potential non-linearity and sparsity, algorithms including Random Forest, LASSO, Gradient Boosting, and Generalized Additive Model are specified in the superlearner when implementing the TMLE.